A 0.05kg ball rotates in a circle of a raduis 0.5 meters.if the ball is travelling at a constant speed of 6.28m/s, find the tention force at the top and bottom of the circle draw a digram for both situations 


1
Expert's answer
2021-06-21T11:40:51-0400


Top:

"m\frac {v^2}{R}=-mg-T,\\\space\\
T=m\bigg(g+\frac{v^2}{g}\bigg)=4.43\text N."

Bottom:


"-m\frac {v^2}{R}=-mg+T,\\\space\\
T=m\bigg(g-\frac{v^2}{g}\bigg)=3.45\text N."